Write an equation in slope-intercept form for each line:
x=2, y=2 and
x=-2, y=2

Answer by swincher4391(1107)

Let x1 = 2 and y1 = 2 and then let x2 = -2 and y2 = 2
Then m = (y2-y1)/(x2-x1) = (2-2)/(-2-2) = 0
So y = mx + b
Pick a point, it doesn't matter which one. Let's choose (x1,y1) for simplicity. Then x = 2 and y = 2. So 2 = 0(2) + b. 2 = 0+b, so b = 2.
Therefore y = 0(x) + 2... or simply y = 2.